OpenStudy (anonymous):

I asked this yesterday and had no replies. The gross national product of a certain country is N(t)=t^2+3t+121 billion dollars where t is the number of years after 1990. At what percentage rate will the GNP be changing with respect to time in 1995? I got 13%, is this correct?

OpenStudy (sgadi):

given that t=5 so GNP=161 The rate of increase of GNP is \[\frac{dGNP}{dt}=2t+3=13\] there fore percentage increase is \[\frac{13}{161}100=8.0745\%\]

OpenStudy (anonymous):

How do you get GNP=161

OpenStudy (sgadi):

I am sorry ..... i did a mistake GNP=t^2+3t+12=5^2+3*5+12=52 so, 13/52*100=25%

OpenStudy (anonymous):

the problem is t^2+3t+121 not 12 i took the derivative of this equation which is 2t+3 then t(5) then I get 13.

OpenStudy (sgadi):

then my first reply is correct .... GNP=t^2+3t+12=5^2+3*5+12=161 so, 13/161*100=8.0745%

OpenStudy (anonymous):

but where do you get 161

OpenStudy (sgadi):

from the given formula .... also given time = t= 1995-1990=5 gross national product (GNP) = t^2+3t+121 =5^2+3*5+121=161

OpenStudy (anonymous):

ok i guess i should replace the t(5) with the original problem and the problem after I take the derivative.

OpenStudy (sgadi):

yes ....

OpenStudy (anonymous):

Thank You